OpenStudy (anonymous):

(csc x + 1)(csc x - 1)

OpenStudy (noelgreco):

FOIL it and apply trig identity

OpenStudy (anonymous):

so its \[(\csc x)^2-1=cot^22x\]

OpenStudy (noelgreco):

yep

OpenStudy (anonymous):

@itsmylife how did you get that

OpenStudy (anonymous):

trig identity

OpenStudy (anonymous):

how

OpenStudy (anonymous):

can you show how you got that plz

OpenStudy (anonymous):

pythagorean identity

OpenStudy (noelgreco):

There are three Pythagorean trig identities that you simply must memorize.

OpenStudy (anonymous):

go here @pvs285 http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

can you just show the work to answer this question plz?

OpenStudy (anonymous):

\[\left( \csc x+1 \right)\left( \csc x-1 \right)=\csc ^{2}x-1=\cot ^{2}x\]

OpenStudy (anonymous):

\[\sin^2x+\cos^2x=1,\frac{ \sin^2 x}{ \sin^2x }+\frac{ \cos^2x }{ \sin^2x }=\frac{ 1 }{ \sin^2x }, 1+\cot^2x=cosec^2x\]

OpenStudy (noelgreco):

It is against the policies of this site to simply provide answers.

OpenStudy (anonymous):

huh u said it was cot^2

OpenStudy (anonymous):

ik i whant you to explain it to me so i uderstand

OpenStudy (anonymous):

it might be a mistake ;)

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

pls help

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

\[\csc ^{2}x-\cot ^{2}=1,or \csc ^{2}x-1=\cot ^{2}x\] It is an identity.

OpenStudy (anonymous):

so there is no work involved?

OpenStudy (noelgreco):

pvs: it appears you are in over your head. If you have a question such as this, and can't pick up on the idea of a trig identity you might want to slow down and back up.

OpenStudy (ybarrap):

$$ \tt \bf \Large(\csc x + 1)(\csc x -1)\\ \Large=\csc^2x-\csc x +\csc x -1\\ \Large=\csc^2x-1\\\\ \Large=\left(\dfrac{1}{\sin x}\right)^2-1\\ \Large=\dfrac{1}{\sin^2x}-1\\ \Large=\dfrac{1}{\sin^2x}-\dfrac{\sin^2x}{\sin^2x}\\\\ \Large=\dfrac{1-\sin^2x}{\cos^2x}\\\\ \Large=\dfrac{\cos^2x}{\sin^2x}\\\\ \Large=\cot^2x $$ Please let me kow if you have any questions. Good luck.